# Evaluating Image Hallucination in Text-to-Image Generation with Question-Answering

### v5

@@ Abstract
Despite the impressive success of text-to-image (TTI) generation models, existing studies overlook the issue of whether these models accurately convey factual information. In this paper, we focus on the problem of image hallucination, where images created by generation models fail to faithfully depict factual content. To address this, we introduce I-HallA (Image Hallucination evaluation with Question Answering), a novel automated evaluation metric that measures the factuality of generated images through visual question answering (VQA). We also introduce I-HallA v1.0, a curated benchmark dataset for this purpose. As part of this process, we develop a pipeline that generates high-quality question-answer pairs using multiple GPT-4 Omni-based agents, with human judgments to ensure accuracy. Our evaluation protocols measure image hallucination by testing if images from existing text-to-image models can correctly respond to these questions. The I-HallA v1.0 dataset comprises 1.2K diverse image-text pairs across nine categories with 1,000 rigorously curated questions covering various compositional challenges. We evaluate five text-to-image models using I-HallA and reveal that these state-of-the-art models often fail to accurately convey factual information. Moreover, we validate the reliability of our metric by demonstrating a strong Spearman correlation ($\rho$=0.95) with human judgments. We believe our benchmark dataset and metric can serve as a foundation for developing factually accurate text-to-image generation models. Additional resources can be found on our project page: https://github.com/hchoi256/I-HallA-v1.0.
@@ Introduction
As generative models (gm:00; stablediff) continue to evolve, the demand for generating factual content alongside imaginary content has grown (gm:03; reimagen). In natural language generation, outputs with factual errors are classified as hallucinations, and considerable research has focused on mitigating this issue (llmfaithfulness1; factscore).
@@
Current text-to-image (TTI) models also struggle to accurately reflect factual information, generating incorrect images for given prompts, as illustrated in Figure 1. This issue is becoming increasingly critical as TTI models are being actively utilized in industries and fields where factual accuracy is essential (problem_nature). For instance, if images against factual information are used in educational materials or the media, misinformation and misconceptions can spread rapidly, causing serious social side effects (problem_google).
@@
While hallucinations have been primarily discussed in the language domain, relatively little research has addressed this issue in the context of image generation. This paper focuses on the unexplored issue of “image hallucination,” where generated images fail to reflect factual information. To understand image hallucination in TTI models and guide future research directions, well-defined evaluation protocols and benchmark datasets are essential. Recent studies have developed benchmarks and evaluation metrics to assess TTI models based on the alignment between text prompts and generated images (vq2_seetrue). These evaluation protocols, such as TIFA (TIFA), focus only on the elements explicitly mentioned in the text prompt. However, verifying the factual alignment of generated images requires external knowledge beyond the prompt. As shown in Figure 1-(a), the number of hydrogen atoms, though not mentioned in the prompt, is important factual information. Additionally, current metrics struggle to distinguish between images that accurately represent factual information and those that simply match the text prompt, especially when polysemy introduces various interpretations. As shown in Figure 1-(b), while the generated image captures the idea of a “comb pattern,” it also includes several incorrect details, like small circles and decorations, which are not part of the factual image.
@@
To address these limitations, we propose a three-stage pipeline to construct a new benchmark, I-HallA v1.0, with a new evaluation metric, I-HallA. Unlike existing protocols, we leverage the vast knowledge base of GPT-4 Omni (GPT-4o) (gpt4o) to assess factual information not mentioned in the prompt, such as the number of hydrogen and carbon atoms, molecular structure, and more. Additionally, we utilize the visual understanding capabilities of GPT-4o to discern factual visual semantics from potentially false ones, which is difficult to do with text prompts alone.
@@
First, the dataset includes 200 prompts based on content from five science and history textbooks (text_history1; text_history2; text_physics; text_biology; text_earthscience) to address factual information. Textbooks, meticulously edited for educational purposes, represent years of accumulated knowledge and are among the most authoritative sources, making them a primary basis for this dataset. Specifically, we use the textbooks’ captions and corresponding figures as prompts representing factual information and factual images. This is because textbook figures are carefully curated, highly aligned with their captions, and thoroughly validated for factual accuracy. The hallucinated images generated from these prompts in five TTI models (stablediff; sd2.0; SDXL; dalle3) are compared against the factual images. In total, we gather 1,200 images for all prompts, consisting of both factual and hallucinated images.
@@
In language models, hallucination refers to the generation of unfaithful content to the given source material (llmhallucsurvey). As large language models (LLMs) increasingly produce text that closely resembles human writing, there has been a growing emphasis on developing benchmarks to evaluate and distinguish hallucinated content. For instance, FEVER (fever) is a dataset used for fact-checking that utilizes Wikipedia as its knowledge source. HaluEval (halueval) combines automated generation with human annotation to detect hallucinations.
Hallucination also occurs in large vision-language models (VLMs) such as LLaVA (llava), where visual features are input into LLMs to generate textual descriptions. In VLMs, hallucination refers to a mismatch between the factual details of images (e.g., object presence, attributes, spatial relations) and the corresponding generated text (vlmsurvey). Various studies evaluate this using metrics like BLEU (bleu) or CIDEr (cider), or by querying VLMs about object presence (pope). In contrast, we focus on evaluating hallucinations in image generation, where generated images fail to depict factual information accurately. While hallucinations can occur in our pipeline during text generation, we mitigate this issue through rigorous human reviews. Further details are covered in Appendix B.
Common Sense Reasoning in Vision-Language Models
Some studies use benchmark datasets to assess whether VLMs possess commonsense knowledge when interpreting images. For instance, the WHOOPS (whoops) and ROME (rome) datasets are created by inputting intentional text prompts that defy common sense into TTI models, resulting in odd and unconventional images.
@@
CLIPScore (clipscore) and DALL-Eval (dalleval), which are early studies measuring the alignment between generated images and text prompts to evaluate TTI models, commonly exhibit limitations due to the inherent constraints of CLIP (e.g., inability to count objects) or the restricted scope of evaluation criteria.
With the growing capabilities of foundation models, a new approach has developed that validates alignment in text-to-image generation by using VQA models on questions derived from the prompt. For instance, TIFA (TIFA) classifies elements of the text prompt into 12 categories and generates a set of questions and answers using GPT-3 (gpt3). These sets are used to evaluate the image by inputting both the image and questions into the VQA model like mPLUG (mplug). VQ2 (vq2_seetrue) extracts key information from the text prompt, generates related questions, and evaluates the text-image alignment by checking whether the image provides correct answers. VPEval (vpeval) enhances these alignment evaluation methods by incorporating object detection and optical character recognition, allowing for a more precise assessment. Davidsonian Scene Graph (dsg) breaks down the prompt into small propositions and represents the dependencies between these propositions in a graph, ensuring that the generated questions are not redundant. However, existing benchmarks that focus solely on evaluating the alignment between text prompts and images often fail to detect external knowledge beyond the text and the factual visual semantics embedded in the image, which is required to address image hallucination. Details are covered in the Methodology and Appendix D.
@@
Factual information refers to data that can be objectively verified and proven true based on evidence or reliable sources. It is an important evaluation criterion in fields that require reliable and accurate information, such as education (fact_important). In this paper, we focus on image hallucination, a phenomenon where the images generated by TTI models fail to accurately reflect factual information.
@@
As shown in Figure 1, existing evaluation metrics, such as TIFA (TIFA), rely solely on text prompts, which limit their ability to consider factual information not explicitly stated in the prompt. In contrast, our metric leverages GPT-4o’s capability to generate questions and answers based on external factual information that is not specified in the prompt but has already been trained into the model. This allows us to assess whether crucial factual information, though not mentioned in the prompt, is accurately reflected in the generated image.

### v7

@@ Abstract
Despite the impressive success of text-to-image (TTI) generation models, existing studies overlook the issue of whether these models accurately convey factual information. In this paper, we focus on the problem of image hallucination, where images created by generation models fail to faithfully depict factual content. To address this, we introduce I-HallA (Image Hallucination evaluation with Question Answering), a novel automated evaluation metric that measures the factuality of generated images through visual question answering (VQA). We also introduce I-HallA v1.0, a curated benchmark dataset for this purpose. As part of this process, we develop a pipeline that generates high-quality question-answer pairs using multiple GPT-4 Omni-based agents, with human judgments to ensure accuracy. Our evaluation protocols measure image hallucination by testing if images from existing TTI models can correctly respond to these questions. The I-HallA v1.0 dataset comprises 1.2K diverse image-text pairs across nine categories with 1,000 rigorously curated questions covering various compositional challenges. We evaluate five TTI models using I-HallA and reveal that these state-of-the-art models often fail to accurately convey factual information. Moreover, we validate the reliability of our metric by demonstrating a strong Spearman correlation ($\rho$=0.95) with human judgments. We believe our benchmark dataset and metric can serve as a foundation for developing factually accurate TTI generation models. Additional resources can be found on our project page: https://sgt-lim.github.io/I-HallA.
@@ Introduction
As generative models (Reimers and Gurevych 2019; Rombach et al. 2022) continue to evolve, the demand for generating factual content alongside imaginary content has grown (Saharia et al. 2022; Chen et al. 2022). In natural language generation, outputs with factual errors are classified as hallucinations, and considerable research has focused on mitigating this issue (Maynez et al. 2020; Min et al. 2023).
@@
Current text-to-image (TTI) models also struggle to accurately reflect factual information, generating incorrect images for given prompts, as illustrated in Figure 1. This issue is becoming increasingly critical as TTI models are being actively utilized in industries and fields where factual accuracy is essential (Wong 2024). For instance, if images against factual information are used in educational materials or the media, misinformation and misconceptions can spread rapidly, causing serious social side effects (Robertson 2024).
@@
While hallucinations have been primarily discussed in the language domain, relatively little research has addressed this issue in the context of image generation. This paper focuses on the unexplored issue of “image hallucination,” where generated images fail to reflect factual information (Lim and Shim 2024). To understand image hallucination in TTI models and guide future research directions, well-defined evaluation protocols and benchmark datasets are essential. Recent studies have developed benchmarks and evaluation metrics to assess TTI models based on the alignment between text prompts and generated images (Yarom et al. 2024). These evaluation protocols, such as TIFA (Hu et al. 2023), focus only on the elements explicitly mentioned in the text prompt. However, verifying the factual alignment of generated images requires external knowledge beyond the prompt. As shown in Figure 1-(a), the number of hydrogen atoms, though not mentioned in the prompt, is important factual information. Additionally, current metrics struggle to distinguish between images that accurately represent factual information and those that simply match the text prompt, especially when polysemy introduces various interpretations. As shown in Figure 1-(b), while the generated image captures the idea of a “comb pattern,” it also includes several incorrect details, like small circles and decorations, which are not part of the factual image.
@@
To address these limitations, we propose a three-stage pipeline to construct a new benchmark, I-HallA v1.0, with a new evaluation metric, I-HallA. Unlike existing protocols, we leverage the vast knowledge base of GPT-4 Omni (GPT-4o) (OpenAI 2024c) to assess factual information not mentioned in the prompt, such as the number of hydrogen and carbon atoms, molecular structure, and more. Additionally, we utilize the visual understanding capabilities of GPT-4o to discern factual visual semantics from potentially false ones, which is difficult to do with text prompts alone.
@@
First, the dataset includes 200 prompts based on content from five science and history textbooks (Jackson J. Spielvogel 2008; Danzer et al. 2008; Urone et al. 2020; O’Grady et al. 2021; Neser 2023) to address factual information. Textbooks, meticulously edited for educational purposes, represent years of accumulated knowledge and are among the most authoritative sources, making them a primary basis for this dataset. Specifically, we use the textbooks’ captions and corresponding figures as prompts representing factual information and factual images. This is because textbook figures are carefully curated, highly aligned with their captions, and thoroughly validated for factual accuracy. The hallucinated images generated from these prompts in five TTI models (Rombach et al. 2022; StabilityAI 2022; Podell et al. 2023; OpenAI 2023) are compared against the factual images. In total, we gather 1,200 images for all prompts, consisting of both factual and hallucinated images.
@@
In language models, hallucination refers to the generation of unfaithful content to the given source material (Ji et al. 2023). As large language models (LLMs) increasingly produce text that closely resembles human writing, there has been a growing emphasis on developing benchmarks to evaluate and distinguish hallucinated content. For instance, FEVER (Thorne et al. 2018) is a dataset used for fact-checking that utilizes Wikipedia as its knowledge source. HaluEval (Li et al. 2023a) combines automated generation with human annotation to detect hallucinations.
Hallucination also occurs in large vision-language models (VLMs) such as LLaVA (Liu et al. 2024a), where visual features are input into LLMs to generate textual descriptions. In VLMs, hallucination refers to a mismatch between the factual details of images (e.g., object presence, attributes, spatial relations) and the corresponding generated text (Liu et al. 2024b). Various studies evaluate this using metrics like BLEU (Papineni et al. 2002) or CIDEr (Vedantam, Zitnick, and Parikh 2015), or by querying VLMs about object presence (Li et al. 2023b). In contrast, we focus on evaluating hallucinations in image generation, where generated images fail to depict factual information accurately. While hallucinations can occur in our pipeline during text generation, we mitigate this issue through rigorous human reviews. Further details are covered in Appendix B.
Common Sense Reasoning in VLMs
Some studies use benchmark datasets to assess whether VLMs possess commonsense knowledge when interpreting images. For instance, the WHOOPS (Guetta et al. 2023) and ROME (Zhou et al. 2023) datasets are created by inputting intentional text prompts that defy common sense into TTI models, resulting in odd and unconventional images.
@@
CLIPScore (Hessel et al. 2021) and DALL-Eval (Cho, Zala, and Bansal 2023), which are early studies measuring the alignment between generated images and text prompts to evaluate TTI models, commonly exhibit limitations due to the inherent constraints of CLIP (e.g., inability to count objects) or the restricted scope of evaluation criteria.
With the growing capabilities of foundation models, a new approach has developed that validates alignment in text-to-image generation by using VQA models on questions derived from the prompt. For instance, TIFA (Hu et al. 2023) classifies elements of the text prompt into 12 categories and generates a set of questions and answers using GPT-3 (Brown et al. 2020). These sets are used to evaluate the image by inputting both the image and questions into the VQA model like mPLUG (Li et al. 2022). VQ2 (Yarom et al. 2024) extracts key information from the text prompt, generates related questions, and evaluates the text-image alignment by checking whether the image provides correct answers. VPEval (Cho, Zala, and Bansal 2024) enhances these alignment evaluation methods by incorporating object detection and optical character recognition, allowing for a more precise assessment. Davidsonian Scene Graph (Cho et al. 2024) breaks down the prompt into small propositions and represents the dependencies between these propositions in a graph, ensuring that the generated questions are not redundant. However, existing benchmarks that focus solely on evaluating the alignment between text prompts and images often fail to detect external knowledge beyond the text and the factual visual semantics embedded in the image, which is required to address image hallucination.
@@
Factual information refers to data that can be objectively verified and proven true based on evidence or reliable sources. It is an important evaluation criterion in fields that require reliable and accurate information, such as education (Hew et al. 2014). In this paper, we focus on image hallucination, a phenomenon where the images generated by TTI models fail to accurately reflect factual information.
@@
As shown in Figure 1, existing evaluation metrics, such as TIFA (Hu et al. 2023), rely solely on text prompts, which limit their ability to consider factual information not explicitly stated in the prompt. In contrast, our metric leverages GPT-4o’s capability to generate questions and answers based on external factual information that is not specified in the prompt but has already been trained into the model. This allows us to assess whether crucial factual information, though not mentioned in the prompt, is accurately reflected in the generated image.